OpenStudy (anonymous):

a boat on a lake bobs up and down due to ripples on the water surface. the person in the bboat notices that the crests of the ripples are 6m apart and that a ripple raieses the boat ever 3 s. calculate the speed of the ripples

OpenStudy (kropot72):

The wavelength = 6m Frequency = 1/period = 1/3 speed = wavelegth * frequency = 6 * (1/3) ms^-1

OpenStudy (anonymous):

thanks very much

OpenStudy (kropot72):

You're welcome :)